OpenStudy (anonymous):

what s is the answer 2x^2+7x+3 A.(2x+1)(x+3) B.(2x+3)(x+1) C.(2x-3)(x-1) D.(2x-1)(x-3)

OpenStudy (anonymous):

A.

hero (hero):

2x^2+6x+x+3 =2x(x+3)+1(x+3) = (x+3)(2x+1)